If someone anyone can help me with this problem I'd really appreciate it. The prblem is " A retangular garden has a length that is 5 meter longer than half the width. If the perimeter is 30 meters longer than twice the width, length, find the length." I know how to find the perimeter but this is on a study guide for semester finals I really need help. PLEASE help. thanks
confused??

Answer by ptaylor(1894)

If you know how to find the perimeter, then that's a good start because the formula for the perimeter has to be used.
Let W= width
then the length, L=W/2+5
and perimeter, P=2W+30
As you indicated, we know that 2W+2L=P so our equation to solve is:
2W+2(W/2+5)=2W+30 simplifying:
2W+W+10=2W+30 collecting like terms:
2W-2W+W=30-10
W=20 width
L=W/2+5=20/2+5=
L=15 Length
You can check the answers by substituting into the above equations.
Hope this helps. Happy holidays-----ptaylor